# bioSBM: a random graph model to integrate epigenomic data in chromatin structure prediction

### v1

@@ Abstract
The spatial organization of chromatin within the nucleus plays a crucial role in gene expression and genome function. However, the quantitative relationship between this organization and nuclear biochemical processes remains under debate. In this study, we present a graph-based generative model, bioSBM, designed to capture long-range chromatin interaction patterns from Hi-C data and, importantly, simultaneously link these patterns to biochemical features. Applying bioSBM to Hi-C maps of the GM12878 lymphoblastoid cell line, we identified a latent structure of chromatin interactions, revealing 12 distinct communities that strongly align with known biological annotations. Additionally, we infer a linear transformation that maps biochemical observables, such as histone marks, to the parameters of the generative graph model, enabling accurate genome-wide predictions of chromatin contact maps on out-of-sample data, both within the same cell line, and on the completely unseen HCT116 cell line under RAD21 depletion. These findings highlight bioSBM’s potential as a powerful tool for elucidating the relationship between biochemistry and chromatin architecture and predicting long-range genome organization from independent biochemical data.
@@ I Introduction
The advent of Chromatin Conformation Capture (3C) [5], particularly its now popular derivative Hi-C [6], allowed the mapping of chromatin contacts at a genome-wide level. Thanks to the fixation of nuclei with formaldehyde, which preserves information about the spatial proximity of linearly distal DNA loci, Hi-C experiments generate contact frequency maps that display non-trivial interaction motifs. One notable feature of the organization of chromatin unveiled by Hi-C is the segregation of the genome into two classes of domains dubbed A and B compartments, which are characterized by distinct interaction patterns. On top of these connectivity differences, A and B compartments have been shown to correlate with epigenetic marks associated with active (A) and silenced (B) transcriptional states [6]. Despite these correlations, more detailed microscopy studies and information from epigenetics (e.g., histone modifications and preferential binding of transcription factors) [7, 8, 9] suggest that the transcriptional state of the genome is more nuanced and the binary classification into A/B compartments may be excessively oversimplified. While A/B compartments and the sub-compartments defined by Rao et al. [9] do provide valuable correlative links between distinct 3D chromatin interaction patterns and different epigenetic marks, a quantitative understanding of the relationship between 3D organization and underlying nuclear biochemical processes remains elusive.
Several attempts to explain the emergence and spatial organization of compartments have invoked polymer-based models, which aim to simulate chromatin structure by incorporating biophysical properties [10, 11, 12, 13]. While such models provide valuable insights by connecting some biochemical processes to mechanisms of genome folding, their reliance on polymer simulations results in high computational costs, particularly when additional parameters need to be estimated. In contrast, more recent deep learning models scale efficiently and have achieved impressive success in predicting Hi-C contact maps. Early models make predictions using DNA sequence alone, but these predictions do not account for cell-type-specific variability [14, 15, 16]. More recently, deep learning approaches have started to incorporate 1D epigenetic signals [17], improving predictive accuracy across different cellular conditions. However, these models remain largely uninterpretable, making it difficult to connect their predictive power with underlying biological mechanisms.
@@
In recent years, ideas from the field of network or graph theory [18] have emerged as a promising paradigm to study chromatin organization at the mesoscopic level. These methods avoid the computational overheads of microscopic polymer-based models by abstracting chromatin structure as a network of interactions, where DNA loci are treated as nodes and their contacts as edges. Such graph-based approaches not only were successfully used to reveal structural patterns [19, 20] but also provided interpretable insights into the relationship between chromatin architecture and biological function [21, 22, 23].
@@
In this paper, we propose bioSBM, an interpretable network model that directly links chromatin structure with biochemical features. bioSBM is based on the stochastic block model (SBM) [24, 25], a class of generative network models that partition the network into communities based on interaction patterns, making them highly suitable for uncovering latent structures in chromatin interaction maps. bioSBM modulates this community structure by considering biochemical covariates such as histone modifications and transcription factor binding, therefore constructing a quantitative framework to understand the relationship between 3D chromatin organization and biochemical processes. Unlike traditional SBMs, which assign each node to a single community, bioSBM allows for mixed memberships, enabling genomic regions to participate in multiple communities simultaneously, thus capturing the context-dependent nature of chromatin interactions.
@@
In Section II we provide a detailed overview of the vanilla stochastic block model and discuss how it can be adapted to better describe different types of analyzed data. In particular, we provide details of the data we utilize (distance-corrected, or observed-over-expected, Hi-C maps) in Sec. II.1, and our customized version of the SBM in Sec. II.2. Then, in Sec. II.3 we introduce the main features of the inference algorithm specifically developed to compute posterior probabilities for our model, while leaving the mathematical details of its derivation in the Supplemental Material (SM). We present our main results in Sec. III, demonstrating the biological relevance and predictive power of our bioSBM model. Finally, in Sec. IV we discuss our results in the context of chromosome organization and conclude by highlighting, in particular, possible future applications.
@@ II.1 SBM and its generalization
SBMs are a particular class of random graphs. A graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ consists of a set of vertices $\mathcal{V}$, representing entities $1,\ldots,N$ and a set of edges $(i,j)\in\mathcal{E}\subset\mathcal{V}\times\mathcal{V}$, denoting pairwise interactions between these entities. Edges can be binary, indicating the presence or the absence of a link, or they can be weighted, reflecting the strength of interactions. Random graphs [26, 27] denote graphs whose structure is itself a random variable, usually as a result of a random attachment process.
@@
SBMs have their roots in the world of social sciences [28, 29], where they were used to model populations divided into sub-populations or communities. The central idea is that interactions between individuals are influenced by their community, creating a non-trivial structure in the interaction graph. This simple yet powerful idea made SBMs into popular models to study general types of relational data.
@@
Numerous algorithms have been developed to detect community structures in complex networks and to make sense of them [30, 31]. Box II.2 provides a schematic overview of the main ideas behind SBMs and of some of their most common extensions.
@@
In our case, the relational data is derived from Hi-C contact frequency maps. Early Hi-C experiments demonstrated that contact frequency between genomic loci is strongly influenced by linear proximity or genomic distance, with the contact probability exhibiting a power-law decay as a function of this distance [6]. This scaling behavior is understood to arise from fundamental polymer physics mechanisms [32, 33, 34]. Instead of using raw Hi-C data, it is often more insightful to study the so-called observed-over-expected (OE) Hi-C maps. These OE maps are derived as the logarithmic ratio between the actual contact frequencies recorded in Hi-C matrices and the expected contact frequencies based on genomic distances. These maps effectively highlight interaction patterns of chromatin by accounting for and removing the global polymeric effects that contribute to the power-law scaling. Significant interactions can be observed across large genomic scales, sometimes spanning entire chromosomes. To uncover a latent structure in these interaction patterns, we employ a weighted SBM. Namely, to capture the complexity of community structures we use a so-called mixed membership version of the SBM (MMSBM, see Sec. II.2) that allows the same genomic regions to belong to multiple classes.
@@
A key difference between standard relational data and the biological setting we consider is the availability of a wealth of additional data in biology. While Hi-C measures contact patterns between chromosomal regions, a variety of biochemical assays, such as ChIP-seq [35], provide 1D genomic maps of specific epigenomic marks at such regions. In our model, we integrate this ChIP-seq data as a vector of biochemical covariates associated with each node, and which modulates the probability of each node belonging to the different communities.
@@
Formally, bioSBM is a hierarchical Bayesian model [36, 37] over weighted graphs. Graph nodes $i\in\{1,\ldots,N\}$ represent a set of contiguous genomic regions of fixed length. The observed weighted network has adjacency matrix $Y$, with $Y_{ij}$ representing the logarithmic OE Hi-C contact frequency (Sec. II.1) between region $i$ and $j$. The observed weighted network is assumed to be generated according to the latent distributions of group memberships for each node/genomic region, as well as the matrices that specify group-group interaction strengths. Each node $i$ has an associated membership proportions’ vector $\theta_{i}$, where $\theta_{ig}$ denotes the probability of node $i$ belonging to group $g$, allowing nodes to belong to multiple classes and display interactions that are context-dependent.
@@
The group-group interaction strengths are defined by matrices $B$ and $\sigma^{2}$, where $B_{gk}$ and $\sigma^{2}_{gk}$ represent the mean and variance of the strength of interaction between class $g$ and class $k$. For each pair of nodes $(i,j)$, discrete variables $z_{ij}$ and $z_{ji}$ denote the group membership of $i$ when interacting with $j$, and vice-versa.
Box 1: Flavours of Stochastic Block Models (SBM’s) SBM’s are a particular type of generative models used in network theory to describe the structure of networks by dividing nodes into communities or blocks. Each block represents a group of nodes that have a similar pattern of connections. The SBM assumes that the probability of a connection between any two nodes depends only on the blocks to which the nodes belong. This model helps understand the network’s underlying structure and is commonly used for community detection [30, 28, 29]. The generative process defined by an SBM is as follows: Determine the number of communities $K$. Assign each node to one of the $K$ communities, For each pair of nodes, generate an edge with a probability that depends on the communities of the nodes. Specifically, generate a Bernoulli random variable with the parameter of the distribution depending on the colors of the two nodes. The left panel of the figure shows a stochastic block model with $K=3$ communities. In this example, the graph is an assortative SBM, meaning that intra-class edge probabilities are higher than inter-class ones. [Uncaptioned image] In its basic version, the SBM is binary, i.e. the edges between nodes are either present or absent. However, many real-world networks involve weighted edges, where the connections between nodes have different strengths or capacities. To adapt the binary SBM for valued (weighted) graphs, we can modify the probability distribution of the edges given the colors or communities to which the two involved nodes belong. Instead of a Bernoulli random variable, we might use Poisson random variables for integer value edges, or Gaussian random variables for real-valued networks, specifying the means and variances of the distributions for each pair of distinct communities [38]. The central panel shows an instance of such weighted SBM. Another aspect we can tweak is the fact that in the traditional SBM, each node belongs to a single community or block. In many networks, nodes may exhibit characteristics of multiple communities. The mixed membership SBM (MMSBM) [39] addresses this by allowing nodes to belong to multiple communities by specifying a probability distribution over classes, or membership proportions. In this paper, we will work with an SBM that has real-valued edges and mixed membership proportions. The right panel shows an example of weighted MMSBM.
Figure 1: Generative process of bioSBM. (a) Hierarchical Bayesian model representing the generative process. The colored circles represent observed variables, which are the edge weights $Y_{ij}$ and the node covariates vectors $x_{i}$. The connectivity patterns of the graph are determined by per-node latent class membership vectors $\theta_{i}$ and global hyperparameters represented by the circles outside the rectangles. (b) An example displays the sampling for the weights of edges between a subset of three nodes.
@@
The edge weight is then sampled from a Gaussian distribution parameterized by $B$ and $\sigma^{2}$ matrices. Putting everything together, the generative process for bioSBM procedes as follows:
@@
  • For every node $i$ we sample a distribution $\theta_{i}$ from the logistic normal distribution [40, 41] with mean $\mu_{i}=\Gamma(x_{i})$ and global covariance $\Sigma$
    $\eta_{i}$ $\sim$ $\mathcal{N}_{K}(\Gamma(x_{i}),\Sigma)$ (1)
    $\theta_{ik}$ $=$ $\frac{\exp(\eta_{ik})}{\sum_{k^{\prime}=1}^{K}\exp{\eta_{ik^{\prime}}}}$ (2)
@@
  • For every pair of nodes $(i,j)$ with $i=1,\ldots,N$ and $j=1,\ldots,i-1$
    $z_{ij}\sim\text{Mult}(\theta_{i})\qquad z_{ji}\sim\text{Mult}(\theta_{j})$ (3)
    with $z_{ij}$ being the membership of node $i$ interacting with node $j$ and vice-versa, sampled from the multinomial distribution with probabilities $\theta_{i}$ and $\theta_{j}$ respectively.
  • Once the memberships of $i$ and $j$ are sampled the weight of the edge is sampled from a Gaussian of which parameters are encoded in the global $B$ and $\sigma$ matrices.
    $P(Y_{ij}|z_{ij}=k,z_{ji}=g,B,\sigma)=\mathcal{N}(Y_{ij}|B_{kg},\sigma_{kg})$ (4)
Schematically, the bioSBM model is represented as a graphical model as in Fig. 1.

### v2

@@ Abstract
The spatial organization of chromatin within the nucleus plays a crucial role in gene expression and genome function. However, the quantitative relationship between this organization and nuclear biochemical processes remains under debate. In this study, we present a graph-based generative model, bioSBM, designed to capture long-range chromatin interaction patterns from Hi-C data and, importantly, simultaneously link these patterns to biochemical features. Applying bioSBM to Hi-C maps of the GM12878 lymphoblastoid cell line, we identified a latent structure of chromatin interactions, revealing $7$ distinct communities that strongly align with known biological annotations. Additionally, we infer a linear transformation that maps biochemical observables, such as histone marks, to the parameters of the generative graph model, enabling accurate genome-wide predictions of chromatin contact maps on out-of-sample data, both within the same cell line, and on the completely unseen HCT116 cell line under RAD21 depletion. These findings highlight bioSBM’s potential as a powerful tool for elucidating the relationship between biochemistry and chromatin architecture and predicting long-range genome organization from independent biochemical data.
@@ I Introduction
The advent of Chromatin Conformation Capture (3C) [5], particularly its now popular derivative Hi-C [6], allowed the mapping of chromatin contacts at a genome-wide level. Thanks to the fixation of nuclei with formaldehyde, which preserves information about the spatial proximity of linearly distal DNA loci, Hi-C experiments generate contact frequency maps that display non-trivial interaction motifs. One notable feature of the organization of chromatin unveiled by Hi-C is the segregation of the genome into two classes of domains dubbed A and B compartments, which are characterized by distinct interaction patterns. On top of these connectivity differences, A and B compartments have been shown to correlate with epigenetic marks associated with active (A) and silenced (B) transcriptional states [6]. Despite these correlations, more detailed microscopy studies and information from epigenetics (e.g., histone modifications and preferential binding of transcription factors [7, 8, 9, 10, 11, 12]) suggest that the transcriptional state of the genome is more nuanced and the binary classification into A/B compartments may be excessively oversimplified. In particular, while A/B compartments and the sub-compartments defined by Rao et al. [9](further studied in numerous other works [13, 14, 15])seem to point out the existence of a direct statistical correlation between epigenetic marks and 3D chromatin interaction patterns, there is comparably much less quantitative understanding of the “microscopic” processes at the origin of these correlations. In general, several attempts to explain the emergence and spatial organization of 3D compartments have invoked polymer-based models [16, 17, 18, 19], which aim at rationalizing the observed structure as the consequence of direct, sequence-specific interactions between distal chromatin loci. While such models connecting the biochemistry of epigenetics to genome folding provide valuable insights, their reliance on polymer simulations results in high computational cost, which becomes especially problematic when one tries to scale simulations up to the size of a typical chromosome [19]. In contrast, more recent deep learning models scale efficiently and have achieved impressive success in predicting Hi-C contact maps. Early models make predictions using DNA sequence alone, but these predictions do not account for cell-type-specific variability [20, 21, 22]. More recently, deep learning approaches have started to incorporate 1D epigenetic signals [23], improving predictive accuracy across different cellular conditions. However, these models remain largely uninterpretable, making it difficult to connect their predictive power with underlying biological mechanisms.
@@
In recent years, ideas from the field of network or graph theory [24] have emerged as a promising paradigm to study chromatin organisation at the mesoscopic level. [25] These methods avoid the computational overheads of microscopic polymer-based models by abstracting chromatin structure as a network of interactions, where DNA loci are treated as nodes and their contacts as edges. Such graph-based approaches not only were successfully used to reveal structural patterns [26, 27, 28] but also provided interpretable insights into the relationship between chromatin architecture and biological function [29, 30, 31].
@@
In this paper, we propose bioSBM, an interpretable network model that directly links chromatin structure with biochemical features. bioSBM is based on the stochastic block model (SBM) [32, 33], a class of generative network models that partition the network into communities based on interaction patterns, making them highly suitable for uncovering latent structures in chromatin interaction maps. The first SBM attempt to model long-range chromatin contacts was proposed in 2015 by Cabreros et al. [26]. bioSBM builds on this previous study by modulating this community structure by considering biochemical covariates such as histone modifications and transcription factor binding, therefore constructing a quantitative framework to understand the relationship between 3D chromatin organization and biochemical processes, a problem already studied in previous works such as [34]. Unlike traditional SBM’s, which assign each node to a single community, bioSBM allows for mixed memberships, enabling genomic regions to participate in multiple communities simultaneously, thus capturing the context-dependent nature of chromatin interactions.
@@
The paper is organized as the following: In Section II we provide a detailed overview of the vanilla stochastic block model and discuss how it can be adapted to better describe different types of analyzed data. In particular, we provide details of the data we utilize (distance-corrected, or observed-over-expected, Hi-C maps) in Sec. II.1, and our customized version of the SBM in Sec. II.2. Then, in Sec. II.3 we introduce the main features of the inference algorithm specifically developed to compute posterior probabilities for our model, while leaving the mathematical details of its derivation in the Supplemental Material (SM [35]). We present our main results in Sec. III, demonstrating the biological relevance and predictive power of our bioSBM model. Finally, in Sec. IV we discuss our results in the context of chromosome organization and conclude by highlighting, in particular, possible future applications.
@@ II.1 SBM and its generalization
SBM’s are a particular class of random graphs. A graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ consists of a set of vertices $\mathcal{V}$, representing entities $1,\ldots,N$ and a set of edges $(i,j)\in\mathcal{E}\subset\mathcal{V}\times\mathcal{V}$, denoting pairwise interactions between these entities. Edges can be binary, indicating the presence or the absence of a link, or they can be weighted, reflecting the strength of interactions. Random graphs [36, 37] denote graphs whose edges (or edge values) are generated according to a probability distribution, making the graph structure itself random.
@@
SBM’s have their roots in the world of social sciences [38, 39], where they were used to model populations divided into sub-populations or communities. The central idea is that interactions between individuals are influenced by their community, creating a non-trivial structure in the interaction graph. This simple yet powerful idea made block models into popular models to study general types of relational data.
@@
Numerous algorithms have been developed to detect community structures in complex networks and to make sense of them [40, 22, 23, 26]. Box II.1 provides a schematic overview of the main ideas behind SBM’s and of some of their most common extensions.
@@
In our case, the relational data is derived from Hi-C contact frequency maps. Early Hi-C experiments demonstrated that contact frequency between genomic loci is strongly influenced by linear proximity or genomic distance, with the contact probability exhibiting a power-law decay as a function of this distance [6]. This scaling behavior can be explained based on fundamental polymer physics mechanisms that shape the three-dimensional folding of chromosomes [41, 42, 43, 44, 45, 46]. Instead of using raw Hi-C data, it is often more insightful to study the so-called observed-over-expected (OE) Hi-C maps. These OE maps are derived as the logarithmic ratio between the actual contact frequencies recorded in Hi-C matrices and the expected contact frequencies based on genomic distances. These maps effectively highlight interaction patterns of chromatin by accounting for and removing the global polymeric effects that contribute to the power-law scaling. Significant interactions can be observed across large genomic scales, sometimes spanning entire chromosomes. To uncover a latent structure in these interaction patterns, we employ a weighted SBM, and to capture the complexity of community structures we use a so-called mixed membership version of the SBM (MMSBM, see Sec. II.2) that allows the same genomic regions to belong to multiple communities.
Box 1: Flavours of Stochastic Block Models (SBM’s) SBM’s are a particular type of generative models used in network theory to describe the structure of networks by dividing nodes into communities or blocks. Each block represents a group of nodes that have a similar pattern of connections. The SBM assumes that the probability of a connection between any two nodes depends only on the blocks to which the nodes belong. This model helps understand the network’s underlying structure and is commonly used for community detection [40, 38, 39]. The generative process defined by an SBM is as follows: Determine the communities and their total number, $K$; Assign each node to one of the $K$ communities; For each pair of nodes, generate an edge with a probability that depends on the communities of the nodes. Specifically, generate a Bernoulli random variable with the parameter of the distribution depending on the colors of the two nodes. The left panel of the figure shows a stochastic block model with $K=3$ communities. In this example, the graph is an assortative SBM, meaning that intra-community edge probabilities are higher than inter-community ones. [Uncaptioned image] In its basic version, the SBM is binary, i.e. the edges between nodes are either present or absent. However, many real-world networks involve weighted edges, where the connections between nodes have different strengths or capacities. To adapt the binary SBM for valued (weighted) graphs, we can modify the probability distribution of the edges given the colors or communities to which the two involved nodes belong. Instead of a Bernoulli random variable, we might use Poisson random variables for integer-value edges, or Gaussian random variables for real-value edges by specifying the means and variances of the distributions for each pair of distinct communities [47]. The central panel shows an instance of weighted SBM with Gaussian edges. Another aspect we can tweak is the fact that in the traditional SBM, each node belongs to a single community or block. In many networks, nodes may exhibit characteristics of multiple communities. The mixed membership SBM (MMSBM) [48] addresses this by allowing nodes to belong to multiple communities by specifying a probability distribution over communities, or membership proportions. In this paper, we will work with an SBM that has real-value edges and nodes with mixed membership proportions. The right panel shows an example of weighted MMSBM.
@@
Figure 1: (a) Forward model of the bioSBM model. Biochemical covariates associated with each node or genomic region are mapped into membership proportions. The membership proportions, together with the interaction strengths between pairs of communities (encoded in the block matrix $\mathbf{B}$), determine the connectivity patterns of the graph. (b) Schematic representation of the inverse inference problem. Given a set of contact maps and covariates matrices, our inference procedure generates estimates of the latent membership proportion for the studied genomic loci, and infers the interaction parameters that characterize the model. (c) the model parameters inferred on training data, is then used to make de novo predictions of contact maps, from biochemical features, for unseen data.
@@
A key difference between standard relational data and the biological setting we consider is the availability of a wealth of additional data in biology. While Hi-C measures contact patterns between chromosomal regions, a variety of biochemical assays, such as ChIP-seq [49], provide 1D genomic maps of specific epigenomic marks at such regions. In our model, we integrate the information from ChIP-seq data as a vector of biochemical covariates associated with each node, and which modulates the probability of each node belonging to the different communities.
@@
Formally, bioSBM is a hierarchical Bayesian model [50, 51] over weighted graphs. Graph nodes $i\in\{1,\ldots,N\}$ represent a set of contiguous genomic regions of fixed length. The observed weighted network has adjacency matrix $Y$, with $Y_{ij}$ representing the logarithmic OE Hi-C contact frequency (Sec. II.1) between region $i$ and $j$. The observed weighted network is assumed to be generated according to the latent distributions of group memberships for each node/genomic region, as well as the matrices that specify group-group interaction strengths. Each node $i$ has an associated membership proportions’ vector $\theta_{i}$, where $\theta_{ig}$ denotes the probability of node $i$ belonging to group $g$, allowing nodes to belong to multiple communities and display interactions that are context-dependent.
@@
The group-group interaction strengths are defined by matrices $B$ and $\sigma^{2}$, where $B_{kg}$ and $\sigma^{2}_{kg}$ represent, respectively, the mean and variance of the strength of interaction between community $k$ and community $g$. For each pair of nodes $(i,j)$, discrete variables $z_{ij}$ and $z_{ji}$ denote the group membership of $i$ when interacting with $j$, and viceversa.
@@
Then, the edge weight is sampled from a Gaussian distribution parameterized by $B$ and $\sigma^{2}$ matrices. Putting everything together, the generative process for bioSBM procedes as follows:
@@
  • For every node $i$, we sample the distribution $\theta_{i}$ from the logistic normal distribution [52, 53] with mean $\mu_{i}=\Gamma(x_{i})$ and global covariance $\Sigma$, i.e.
    $\eta_{i}$ $\sim$ ${\mathcal{N}}(\Gamma(x_{i}),\Sigma)\,$, (1)
    $\theta_{ik}$ $=$ $\frac{\exp(\eta_{ik})}{\sum_{k^{\prime}=1}^{K}\exp(\eta_{ik^{\prime}})}\,.$ (2)
@@
  • For every pair of nodes $(i,j)$ with $i=1,\ldots,N$ and $j=1,\ldots,i-1$, we define:
    $z_{ij}\sim\text{Mult}(\theta_{i})\,,\qquad z_{ji}\sim\text{Mult}(\theta_{j})\,$, (3)
    with $z_{ij}$ being the membership of node $i$ interacting with node $j$ and viceversa, sampled from the multinomial distribution with probabilities $\theta_{i}$ and $\theta_{j}$ respectively.
  • Once the memberships of $i$ and $j$ are sampled, the weight $Y_{ij}$ of the corresponding edge is sampled from a Gaussian whose parameters are encoded in the global $B$ and $\sigma^{2}$ matrices:
    $P(Y_{ij}|z_{ij}=k,z_{ji}=g,B,\sigma^{2})=\mathcal{N}(Y_{ij}|B_{kg},\sigma^{2}_{kg})\,$, (4)
where the notation (used throughout the whole text) $P(\cdot|\cdot)$” stands for the conditional probability of the variable(s) on the left given the variable(s) on the right. Schematically, the bioSBM model is illustrated as a graphical model in Fig. 1.